OpenStudy (anonymous):

a firm can produce at most 100 units each week. if its total cost function is known to be C(x)= 500 + 1500x with a total revenue of R (x) = 1600x + x^2, how many units, x, should the firm produce to maximize its profit? Find the max. profit

OpenStudy (owlfred):

Hoot! You just asked your first question! Hang tight while I find people to answer it for you. You can thank people who give you good answers by clicking the 'Good Answer' button on the right!

OpenStudy (anonymous):

profit = revenue - cost = \[1600x+x^2-(500+1500x) \]

OpenStudy (anonymous):

= \[x^2-1500x+1100\] if my algebra is right yes?

OpenStudy (anonymous):

oh it is R (x)= 1600x - x^2

OpenStudy (anonymous):

well then it makes more sense

OpenStudy (anonymous):

so we get \[-x^2-1500x+1100\] yes?

OpenStudy (anonymous):

wait my algebra is wrong

OpenStudy (anonymous):

\[1600x-x^2-(500+1500x)\]

OpenStudy (anonymous):

this time i get \[-x^2+100x-500\] and now i think it is right. sorry

OpenStudy (anonymous):

maximum is at \[x=-\frac{b}{2a}=-\frac{100}{-2}=50\]

OpenStudy (anonymous):

so firm should produce 50 units and its profit will be whatever you get when you replace x by 50

OpenStudy (anonymous):

yeah sorry my bad 50 into original equation?

OpenStudy (anonymous):

yes. i get 2000

OpenStudy (anonymous):

\[-50^2+100\times 50 -500=2000\]

OpenStudy (anonymous):

but you should check it because i used a calculator

OpenStudy (anonymous):

profit function = 100x + x^2 - 500. differentiate this equation with respect to x and set equal to zero. Then solve for x. you want ur marginal profit from selling an extra unit to be zero.

OpenStudy (anonymous):

profit function is 100x -x^2 - 500. sorry

OpenStudy (anonymous):

differentiating is not necessary. vertex of a quadratic is \[-\frac{b}{2a}\] whether you a) take the derivative, set = 0 and solve, b) complete the square, or c) remember how to get a vertex

OpenStudy (anonymous):

certainly do not need calc for this problem

OpenStudy (anonymous):

yeah i got 2000 so max profit?

OpenStudy (anonymous):

from an economic perspective it makes more sense to want marginal profit zero. but either way obviously works

OpenStudy (anonymous):

so 2000 is max profit?

OpenStudy (anonymous):

yes. produce 50, get a profit of 2000